LASER EXCITED METASTABLE SPECTROSCOPY OF C_{2}H_{2}

\underline{KEVIN L. CUNNINGHAM}, STEPHEN DRUCKER, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge MA 02139-4307; SUSAN J. HUMPHREY, CHRISTOPHER G. MORGAN, and ALEC M. WODTKE, Department of Chemistry, University of California at Santa Barbara, Santa Barbara CA 93106.

Triplet states have been detected near the 3ν_{3} vibrational level of the S_{1} electronic surface of C_{2}H_{2} using the highly selective technique of Laser Excited Metastable (LEM) spectroscopy. LEM detects metastable molecules through the detection of an Auger electron emitted when the molecule impacts a metal surface. A supersonic molecular beam of C_{2}H_{2} was crossed with a laser beam which excites π^{∗} − π transitions to metastable states. The molecules then travel ∼100 μs before impacting a Au surface. During the collision, an electron from the conduction band of the metal may fall into the open π bonding orbital of C_{2}H_{2}. The electron in the π^{∗} orbital is then ejected and detected. To be detected, the excited state must have a minimum lifetime on the order of 100 μs and must have electronic excitation energy that exceeds the work function of the Au surface (ϕ = 5.1 eV). The LEM and LIF spectra were recorded simultaneously for a number of bands in the Ã-X̃ band system, including 3ν_{3}. LEM and LIF offer complementary information. LIF detects short-lived states which are of predominant singlet character; LEM detects long-lived states which are predominantly triplet. The spectra show a number of interesting features necessary to understanding the mechanism of singlet-triplet coupling in C_{2}H_{2}. Despite the low vibrational density of T_{3} states near 3ν_{3}, the results show that a single vibrational level of the T_{3} surface is strongly coupled to 3ν_{3} and to the much denser manifold of T_{1} and T_{2} vibrational states.